1. Calculate ΔU if 100 kJ of heat energy is absorbed by the system and 30 kJ of work is done on the surroundings.
1
Expert's answer
2021-01-20T03:15:30-0500

As given that work is done on the surroundings and heat is absorbed by the system.

Therefore,

q = 100kJ

w = −30kJ

Now, from first law of thermodynamics,

ΔU = q + w

⇒ ΔU = 100 + (−30) = 70kJ

Hence, ΔU is 70kJ